Mr. Vijay, a retired government servant, is considering investing his money in two proposals. He wants to choose the one that has higher average net present value and lower standard deviations. The relevant data are given below. Can you help him choosing the proposal?

Proposal A

Net present value (NPV) Chance of the possible outcome of NPV

1559 0.30

5662 0.40

9175 0.30


Proposal B

Net present value (NPV) Chance of the possible outcome of NPV

-10050 0.30

5812 0.40

20584 0.30



1
Expert's answer
2022-02-22T17:15:45-0500

Proposal A

Given

X 1559 5662 9175

P(X) 0.30 0.40 0.30

Therefore, the mean;

μ = Σ XiPi = 1559*0.30 + 5662*0.40 + 9175= 5485



The variance σ2 = Σ(x – μ)2 ⋅ P(x)

= 0.30*(1559 - 5485)2 + 0.4*(5662-5485)2 + 0.3*(9175-5485)2

=8721404.4

The standard deviation is the square root of the variance

standard deviation = (8721404.4)1/2 = 2953.20

Therefore, the average net present value for proposal A = 5485, with a standard deviation = 2953.20


Proposal B


Also, given

X -10050 5812 20584

P(X) 0.30 0.40 0.30

Therefore, the mean;

μ = Σ XiPi = -10050*0.30 + 5812*0.40 + 20584= 5485



The variance σ2 = Σ(x – μ)2⋅ P(x)

= 0.30*(-10050- 5485)2 + 0.4*(5812-5485)2 + 0.3*(20584-5485)2

= 140837579.4

The standard deviation is the square root of the variance

standard deviation = (140837579.4)1/2 = 11867.50

The average net present value for proposal B = 5485, with a standard deviation = 11867.50.


Although both proposal A and B have the same average net present value (5485), Mr. Vijay should consider investing in proposal A since it has a lower standard deviation (2953.20) as compared to proposal B which has a higher standard deviation (11867.50)